OpenStudy (anonymous):

Find the difference. 6rs -2rs A. 4 B. 3rs C. 4rs D. 8rs

OpenStudy (anonymous):

4rs

OpenStudy (anonymous):

rs are the apples here!

OpenStudy (anonymous):

6 apples minus 2 apples!

OpenStudy (saifoo.khan):

http://www.apple.com

OpenStudy (anonymous):

:-D

OpenStudy (anonymous):

4rs

OpenStudy (anonymous):

I wish I had 4 apple mac books

OpenStudy (anonymous):

6rs-4rs=rs(6-2)=rs(4) = 4rs

OpenStudy (saifoo.khan):

i wanted 5.

OpenStudy (anonymous):

4rs